From: NASA HQ/SpaceRef/NASA Watch
Posted: Monday, November 1, 2004
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.† Underway: Week 2 of Increment 10.
FE Sharipov's station inspection this morning included the regular monthly routine checkup on the DC1 "Pirs" docking module's AZS circuit breakers on the BVP Amp Switch Panel (they should all be On) and the LEDs (light-emitting diodes) of 14 fuses in Fuse Panels BPP-30 & BPP-36.
CDR/SO Chiao conducted a new BCAT-3 (Binary Colloid Alloy Test 3) science photography session.† Prior to initiating BCAT-3 ops, Leroy had some familiarization time scheduled to review the procedures before the MWA (Maintenance Work Area) and experiment setup.†† [After setting up the SGSM (slow growth sample module) in the Lab MWA at EXPRESS Rack 5 (ER5), the Science Officer conducted the video-recorded activity by photographing all ten samples with the Kodak 760 digital still camera, prior to homogenization of Sample 1.† After its homogenization, the sample was then photographed at 100, 1,000, and 10,000 sec intervals.† Photography of Sample 1 at the 100,000 sec interval is scheduled for tomorrow, prior to homogenization and photography of Sample 2.† The sample module was mounted once and will remain on the MWA until after the 100,000 sec interval for Sample 2 on 11/3 (Wednesday).† The images captured during the photo sessions over the next couple of days are to be separated onto three PCMCIA memory cards.]
As part of regular Russian segment (RS) fire alarm maintenance, Salizhan worked in the FGB to clean the needles of the IDZ-2 smoke detectors (except for SD #10, which is powered off).
Chiao conducted another installment of the extensive periodic noise level measurements program throughout the station, using the joint US/Russian sound level meter (SLM) in the cabin for a 2-hr. acoustic survey.† The recorded data were later transferred to the medical equipment computer (MEC).†† [These acoustic measurements are obtained once every two months at 46 locations in the Lab (13), Node (4), Airlock (3), FGB (7), SM (11) and DC-1 (3) modules.† The survey also includes five crew preference locations taken at their perceived loudest locations in the station.† The SLM gives instantaneous noise levels and their frequency spectra, which are transferred to the MEC laptop via an RS232 cable and later downlinked with regular CHeCS (crew health care systems) data dump or via OCA. ]
Salizhan Sharipov set up and worked a proficiency training session on the ADUM (Advanced Diagnostic Ultrasound in Microgravity) experiment, using the ADUM OPE (On-board Proficiency Enhancer) in preparation for the crew's ultrasound scanning activities on Thursday (11/4).†† [The FE used the ADUM OPE compact disk on the HRF PC/laptop, focusing on cardiac, thoracic & bone scanning, plus data acquisition (probe positioning) and principles of remote guidance, ultrasound, and anatomy.]
CDR Chiao conducted an audit/inventory of flight crew equipment (FCE) contained in two completely full half-CTBs (crew transfer bags).†† [To ascertain the contents of the containers including any broken/damaged hardware to be trashed, Leroy opened each bag and reported on its items and the actual time spent on the audit via OCA email message.]
The FE performed the daily routine maintenance of the SOZh environmental control & life support systems in the SM and prepared the daily "delta" file for updating the IMS (inventory management system) databases.
Both crewmembers completed their daily 2.5-hr. physical exercise program on TVIS treadmill, CEVIS, RED exerciser and VELO cycle with NS-1 bungee cord load trainer.
Afterwards, Leroy transferred TVIS and RED exercise files to the MEC.
Working off the crew-discretionary Russian task list, FE Sharipov completed the periodic inspection of the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.
Both crewmembers again had one hour set aside on today's schedule for ISS familiarization, i.e., to adjust to their new surroundings and activities.†† [This "free" session has become a valuable standard requirement for new station occupants for the first two weeks.]
An attempt by the crew to "ghost" another PCS (portable computer system) hard disk drive (HDD) last Saturday failed.† The HDD, #6029, which had a previous failure signature, could not be recovered, and the crew used another unit, #6095, instead, thereby increasing the number of operational spare IBM 760XD PCS laptops to two.†† [If #6092 turns out to be nonrestorable, we will have two totally failed HDDs on orbit.]
P6 Battery Reconditioning Status:† Reconditioning of the P6 truss' 4B solar array's battery #2 (4B2) was successfully completed early this morning.† The 4B2 set of three batteries is back online and the 4B channel is in a nominal configuration, providing power to the vehicle.† New software PPLs (pre-positioned loads) are now being built that compensate for what has been learned about the aging affects on the batteries.† Once the new PPLs have been uploaded to the 4B and 2B MDMs (multiplexer/demultiplexers, =computers), a final discharge/charge cycle will be performed to test the capacity of the reconditioned battery set.† Afterwards BCDU 4B2 will again be brought back on-line.
At ~2:00pm EST, a ground-only conferencing setup with POIC (Payload Operations & Integration Center) via OCA will be tested, using simulated space-to-ground (S/G) on the ISS OCA loop.†† [The U.S. OCA (orbital communication adapter) system provides two-way comm between onboard SSC NGLs (station support computer/next generation laptops) via the onboard OCA router plus TDRS satellites and the ground network, which consists of the main server (CDSS) and five OCA routers of the OCA LAN (local area network) to the various users such as Capcoms, biomed and payload operators as well as the IP phone.† The tests are to verify basic point-to-point ops and Capcom-to-POIC interfaces, with the OCA Officer standing in for the crew.]
Upcoming Key Events:† The current station attitude of sun-pointing XPOP (x-axis perpendicular to orbit plane) will be maintained until 11/16.† Station reboost will then follow on 11/17, with backup opportunity on 11/18.† The Soyuz relocation is scheduled for 11/29 (no earlier than), preceded by a hot-fire test of its thrusters. The new SIGI (space integrated GPS/inertial navigation system) R2 firmware upload activities are GO pending completion of safety procedural modifications.]
No CEO (crew earth observations) targets uplinked today.
CEO images can be viewed at these websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 10 crew visit:
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Altitude History
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //